# Robust inference of gravitational wave source parameters in the presence of noise transients using normalizing flows

### v1

@@ Abstract
Gravitational wave (GW) detection is of paramount importance in fundamental physics and GW astronomy, yet it presents formidable challenges. One significant challenge is the removal of noise transient artifacts known as “glitches,” which greatly impact the search and identification of GWs. Recent research has achieved remarkable results in data denoising, often using effective modeling methods to remove glitches. However, for glitches from uncertain or unknown sources, current methods cannot completely eliminate them from the GW signal. In this work, we leverage the inherent robustness of machine learning to obtain reliable posterior parameter distributions directly from GW data contaminated by glitches. Our network model provides reasonable and rapid parameter inference even in the presence of glitches, without needing to remove them. We also investigate various factors affecting the rationality of parameter inference in our normalizing flow network, including glitch and GW parameters. The results demonstrate that the normalizing flow can reasonably infer the source parameters of GWs even with unknown contamination. We find that the nature of the glitch itself is the only factor that can affect the rationality of the inferred results. With improvements to our model, we anticipate accelerating the localization of electromagnetic counterparts and providing priors for more accurate deglitching, thereby speeding up subsequent data processing procedures.
@@ I Introduction
Approximately a hundred years ago, Einstein predicted the existence of gravitational waves (GWs) within the framework of general relativity. It wasn’t until 2015 that the Advanced LIGO and Virgo GW detectors achieved the first direct detection of a GW signal from the merger of a binary black hole (BBH) system, known as GW150914 Abbott et al. (2016a). This discovery marked a monumental advance in the field.
@@
GW research is pivotal for testing general relativity Berti et al. (2015); Abbott et al. (2016b, 2019a, 2019b, 2021a, 2021b); Gong et al. (2022, 2023), exploring astrophysical and cosmological phenomena, and accurately measuring cosmological parameters Abbott et al. (2017); Chen et al. (2018); Abbott et al. (2021c, 2023a); Soares-Santos et al. (2019); Palmese et al. (2020); Sakstein and Jain (2017); Wang et al. (2018); Zhang et al. (2019a); Zhang (2019); Zhang et al. (2020); Wang et al. (2020a); Zhang et al. (2019b); Zhao et al. (2020); Verde et al. (2019); Palmese et al. (2023); Jin et al. (2020); Wang et al. (2022); Jin et al. (2021); Qi et al. (2021); Song et al. (2024); Jin et al. (2022); Wu et al. (2023); Jin et al. (2023a); Li et al. (2024); Jin et al. (2024); Han et al. (2023); Jin et al. (2023b); Dong et al. (2024); Feng et al. (2024). These activities significantly enhance our understanding of fundamental physics and the evolution of the universe.
@@
However, the data received by detectors is highly susceptible to interference from noise transient, resulting in deviation in the analysis process Abbott et al. (2019c, 2021d, 2023b, 2018a). The LIGO-Virgo-KAGRA collaboration has revealed that non-Gaussian features in the data, such as noise transient, or called “glitch” Abbott et al. (2020); Davis et al. (2021); Acernese et al. (2023), can significantly impact the detector’s ability to search for GW signals Abbott et al. (2020).
@@
Current observations have characterized dozens of different glitches by Gravity Spy Zevin et al. (2017); Soni et al. (2021). These glitches can exhibit variations in terms of their frequency, amplitude, and duration, making their identification and mitigation complex tasks. They are categorized according to time, frequency, signal-to-noise ratio (SNR), and duration Robinet et al. (2020); Lynch et al. (2017), with the most common being blip glitches, scattered-light glitches and thunderstorm glitches Razzano and Cuoco (2018); Zevin et al. (2024); Hourihane et al. (2022). Indeed, glitches have significant impacts on GW data and subsequent data analysis Abbott et al. (2016c); Aasi et al. (2014); Abbott et al. (2016d). These non-Gaussian non-stationary features can mimic or mask the signature of GWs, making it more difficult to distinguish real signals from noise Nitz (2018); Cabero et al. (2019). For example, they can affect the search phase, causing detectors to generate false alarms Abbott et al. (2018a). The presence of glitches causes incorrect sky positioning in GW data analysis, affects the tracking of electromagnetic (EM) counterparts Pankow et al. (2018), and biases the extrapolation of cosmological parameters Macas et al. (2022); Zhu and Chen (2023).
@@
The interference caused by glitches in GW data cannot be ignored, and many efforts are made to mitigate their disturbance Cornish (2021); Powell (2018); Steltner et al. (2022); Talbot et al. (2021); Soni et al. (2020); Hourihane et al. (2022); Mohanty and Chowdhury (2023); Udall and Davis (2023). The gating approach uses a windowing function to directly smooth the portion affected by the glitch for events whose time-frequency can be separated Usman et al. (2016). Gwsubtract Davis et al. (2019, 2022); Abbott et al. (2023b) relies on the auxiliary channel to remove glitches, but it has a limited impact on removing glitches that do not appear in auxiliary channels. BayesWave reconstructs the Gaussian noise, glitches, and GW signals using wavelet transforms and inter-correlations between detectors to remove glitches Cornish and Littenberg (2015); Cornish et al. (2021); Littenberg et al. (2016). However, this high-dimensional Bayesian analysis is extremely computationally intensive, and individual events cannot be analyzed in a timely and efficient manner at expected future event rates. Therefore, the development of an effective framework for rapidly mitigating the impacts of glitches and inferring GW parameters has become an urgent need.
@@
The development of Artificial Intelligence (AI) offers an excellent way to facilitate GW research by overcoming the problems of high computational costs and handling huge amounts of data. Currently, most state-of-the-art machine learning algorithms remove non-Gaussian noise components by fitting their parameters Biswas et al. (2013); Yu and Adhikari (2022); Vajente et al. (2020); Ormiston et al. (2020); Merritt et al. (2021); Ashton (2023); Bini et al. (2023). Moreover, machine learning techniques utilize the normalizing flows for the posterior distribution of source parameters Dax et al. (2021); Du et al. (2024); Green and Gair (2021); Williams et al. (2021); Langendorff et al. (2023); Ruhe et al. (2022); Ruan et al. (2023a); Wang et al. (2020b); Ruan et al. (2023a); Zhao et al. (2023); Ruan et al. (2023b); Ma et al. (2024); Gabbard et al. (2018); Fan et al. (2019); Yun et al. (2023a, b); Zhao et al. (2024); Dax et al. (2023). While these methods can effectively enhance data quality in certain aspects, they cannot ensure that the noise remaining after deglitching consists solely of Gaussian noise. Recently, alternative approaches have not involved the removal or subtraction of non-Gaussian noise Du et al. (2024); Sun and Li (2023). Instead, these approaches focus on incorporating non-Gaussian noise into the training process and utilizing likelihood-free methods to directly infer GW parameters Gabbard et al. (2022).
@@
In our previous work, we employed the normalizing flow to perform parameter inference on GW signals contaminated by several common glitches Sun et al. (2024). However, glitches come in different types with varying frequencies of occurrence, and not all types of glitches can be well modeled or have sufficient numbers for training. We are seeking a method that does not rely on modeling to remove glitches and enable rapid parameter inference. Recent studies have demonstrated the robustness of machine learning Wei and Huerta (2020); Wang et al. (2024, 2023); Chatterjee et al. (2021); Xu et al. (2024). The robustness of machine learning methods can aid in performing parameter inference reasonably, even without explicitly modeling and removing glitches. It enables the effective mitigation of non-Gaussian noise interference while inferring GW parameters.
@@
In this work, our primary aim is to investigate the impact of noise transients on the parameter inference obtained from the normalizing flow network. We can use Gaussian noise for training to infer signal parameters contaminated by non-Gaussian noise. Specifically, we utilize simulated sine-Gaussian glitches, encompassing a wide range of morphologies found in advanced detector noise transients Powell et al. (2015); Bose et al. (2016). We discuss the effect of the number of detectors on robust inference in the normalizing flow network. We find that our model provides a relatively better posterior distribution compared to the Bayesian inference method when dealing with glitch-contaminated GW data. In this work, we utilize the bilby dynesty Ashton et al. (2019); Speagle (2020) to implement Bayesian inference. We have also explored the impact of the GW signal intensity on the robustness of network inference. The results indicate that the SNR of the GW signal has varying degrees of impact on the network’s ability to perform source parameter inference. Furthermore, we consider the influence of glitches’ intensity on the robustness of source parameter inference when utilizing the normalizing flow network. Meanwhile, we conduct a detailed study on the impact of three key features of non-stationary noise, i.e., frequency $f$, quality factor $Q$, and amplitude $h_{0}$ Canton et al. (2014), on the robustness inference using the normalizing flow. We find that it is possible to determine the feasibility of obtaining reasonable results by assessing the robustness of the inference considering the characteristics of non-Gaussian noise.
@@
The results demonstrate that the normalizing flow network can provide reliable posterior parameter inference for GWs contaminated by unknown glitches without undergoing deglitching. When GW data is contaminated by unknown noise, our model is capable of delivering swift and reasonable parameter inference, albeit with a difference in magnitude. This provides a reasonable prior for subsequent more accurate processing. When extended to binary neutron star mergers (BNS), this method is also expected to enable the rapid localization of EM counterparts.
@@
This paper is structured as follows. In Section II the methodology used in the work is described in detail. Some of the work’s conclusions and discussions are presented in Section III. Finally, we present the main conclusions of our work in Section IV.
@@ II.1 Normalizing flow network architecture
The normalizing flow is a deep learning-based generative model for learning complex probability distributions Rezende and Mohamed (2015); Durkan et al. (2019). Its core idea is to transform a simple probability distribution into a complex probability distribution through a series of reversible transformations. In the inference process, the normalizing flow generates samples by applying a series of reversible transformations to the base distribution. The normalizing flow network learns the parameters of these transformations to map the base distribution to a posterior distribution corresponding to the real data distribution.
The normalizing flow also has a wide range of applications in generative modeling, where they can be used for tasks such as sample generation, probability density function calculation, and data reconstruction. As the normalizing flow is reversible, both the generation and inference processes are solvable, making it a powerful and flexible modeling technique.
The normalizing flow Rezende and Mohamed (2015); Kingma et al. (2016); Papamakarios et al. (2017) can be expressed by the following formula
@@
Our architecture for normalizing flow is quite similar to the one described in Ref. Sun et al. (2024), except for data fusion. The normalizing flow network consists of a series of coupling transformations, each of which is transformed according to the input data and the output of the previous layer of the neural network, allowing the predicted distribution to gradually approximate the posterior distribution. Between coupling transformations, the order of the sample components is randomized to ensure that the transformation sequence specifically fully transforms all components, we use the rational quadratic spline coupling transformation. Unlike the parameters used in Ref. Sun et al. (2024), the network parameters in this work are shown in Table 1. For networks corresponding to different numbers of detectors, their network parameters remain the same, except that the number of channels in the network corresponds to the number of detectors.
The overall structure of the neural network is illustrated in Fig. 1. The data strain input to the network first undergoes the ResNet neural network, then the feature vectors extracted from the ResNet model. Finally, the feature vectors are processed by the normalizing flow to obtain the GW waveform as well as the parameter inference.
Figure 1: The workflow of the normalizing flow network. The input strain data contains simulated GW signals and Gaussian noise. The strain data is passed through an identical residual net (ResNet) model, and then the feature vectors are extracted from the ResNet model. Then it proceeds through the normalizing flow to obtain the parameter posterior distributions. The above steps are repeated by injecting glitches into the strain data as a test set.
@@ II.2 Data set
In terms of dataset preparation, the data used during the training phase consists of simulated Gaussian noise and simulated GW signals. The signals used are from the GW signals simulated by the SEOBNRv4_opt model Bohé et al. (2017), and they are selected in the range of SNR between 8 and 32. The noise used is a simulated Gaussian noise generated using Pycbc Biwer et al. (2019); Dal Canton et al. (2014) with a frequency range of $20-4096$ Hz, which represents the high-pass frequency and sampling frequency respectively. To obtain the noise sensitivity curves ${S_{n,k}(f)}$ for different detectors, we utilize the aLIGOZeroDetHighPower, AdvVirgo, and KAGRA methods from the pycbc.psd library Nitz and othes (2022); Sellers et al. (2022). After generating Gaussian noise, we inject a signal between 0.7 s and 0.9 s. The specific parameter distribution of the signal is shown in Table 2, and the simulated GW observation data is whitened. Finally, a 2-second dataset is generated.
@@
where $\bm{\tilde{h}}$ is the frequency domain GW waveform considering the detector network including $N$ independent detectors and can be written as $\bm{\tilde{h}}=\left[\tilde{h}_{1},\tilde{h}_{2},\cdot\cdot\cdot,\tilde{h}_{k},\cdot\cdot\cdot,\tilde{h}_{N}\right]$. Here $\tilde{h}_{k}$ is the frequency domain GW waveform of the k-th detector. The inner product is defined as Jin et al. (2024)
@@
and ${S_{n,k}(f)}$ is the sensitivity curve function of the k-th detector Owen and Sathyaprakash (1999). In the above expression, $f_{\rm min}$ is the high-pass frequency, and $f_{\rm max}$ is the Nyquist frequency. Also, we calculate the glitch SNR in the same way. To ensure the validity of the normalizing flow network predictions and prevent overfitting, we regenerate 10,000 new data points at the beginning of each epoch and employ 50 cycles of early stopping.
@@
In the test dataset, we utilize simulated Gaussian noise and simulated GW signals, injecting simulated glitches into them, as illustrated in Fig. 2. The glitches we use are modeled using sine-Gaussian functions Blackburn et al. (2008), which cover most of the types of glitches currently observed by the LIGO, Virgo, and KAGRA networks. We examine a sine-Gaussian glitch Bose et al. (2016), which is defined as
@@
where $\tau=Q/\sqrt{2}\pi f_{0}$ is the time duration. Here $h_{0}$ is the amplitude, $Q$ is the dimensionless quality factor, $t_{0}$ and $f_{0}$ are the location of the sine-Gaussian in time and frequency respectively. We set $t_{0}=0.8$ s to simulate the scenario where the signal and glitch overlap. The selected parameters range of glitches are shown in Table 3.

### v2

# thanks: Corresponding author

@@ Abstract
Gravitational wave (GW) detection is of paramount importance in fundamental physics and GW astronomy, yet it presents formidable challenges. One significant challenge is the removal of noise transient artifacts known as “glitches”, which greatly impact the search and identification of GWs. Recent research has achieved remarkable results in data denoising, often using effective modeling methods to remove glitches. However, for glitches from uncertain or unknown sources, current methods cannot completely eliminate them from the GW signal. In this work, we leverage the inherent robustness of machine learning to obtain reliable posterior parameter distributions directly from GW data contaminated by glitches. Our network model provides reasonable and rapid parameter inference even in the presence of glitches, without needing to remove them. We also investigate various factors affecting the rationality of parameter inference in our normalizing flow network, including glitch and GW parameters. The results demonstrate that the normalizing flow can reasonably infer the source parameters of GWs even with unknown contamination. We find that the nature of the glitch itself is the only factor that can affect the rationality of the inferred results. With improvements to our model, we anticipate accelerating the localization of electromagnetic counterparts and providing priors for more accurate deglitching, thereby speeding up subsequent data processing procedures.
@@ I Introduction
Approximately a hundred years ago, Einstein predicted the existence of gravitational waves (GWs) within the framework of general relativity. It wasn’t until 2015 that the Advanced LIGO detectors achieved the first direct detection of a GW signal from the merger of a binary black hole (BBH) system, known as GW150914 Abbott et al. (2016a). This discovery marked a monumental advance in the field.
@@
GW research is pivotal for testing general relativity Berti et al. (2015); Abbott et al. (2016b, 2019a, 2019b, 2021a, 2021b); Gong et al. (2022, 2023), exploring astrophysical and cosmological phenomena, and accurately measuring cosmological parameters Abbott et al. (2017); Sakstein and Jain (2017); Chen et al. (2018); Soares-Santos et al. (2019); Verde et al. (2019); Jin et al. (2020); Palmese et al. (2020); Abbott et al. (2021c); Jin et al. (2021); Wang et al. (2022); Guo (2022); Jin et al. (2022); Abbott et al. (2023a); Palmese et al. (2023); Song et al. (2024); Li et al. (2024); Jin et al. (2024a); Han et al. (2024); Jin et al. (2024b). These activities significantly enhance our understanding of fundamental physics and the evolution of the universe.
@@
However, the data received by detectors is highly susceptible to interference from noise transient, resulting in deviation in the analysis process Abbott et al. (2018a, 2019c, 2021d, 2023b). The LIGO-Virgo-KAGRA collaboration has revealed that non-Gaussian features in the data, such as noise transient, or called “glitch” Abbott et al. (2020); Davis et al. (2021); Acernese et al. (2023), can significantly impact the detector’s ability to search for GW signals Abbott et al. (2020).
@@
Current observations have characterized dozens of different glitches by Gravity Spy Zevin et al. (2017); Soni et al. (2021). These glitches can exhibit variations in terms of their frequency, amplitude, and duration, making their identification and mitigation complex tasks. They are categorized according to time, frequency, signal-to-noise ratio (SNR), and duration Robinet et al. (2020); Lynch et al. (2017), and there are many kinds of glitches, among which the common glitches being blip glitches, scattered-light glitches and thunderstorm glitches Razzano and Cuoco (2018); Hourihane et al. (2022); Zevin et al. (2024). Indeed, glitches have significant impacts on GW data and subsequent data analysis Aasi et al. (2014); Abbott et al. (2016c, d). These non-Gaussian non-stationary features can mimic or mask the signature of GWs, making it more difficult to distinguish real signals from noise Nitz (2018); Cabero et al. (2019). For example, they can affect the search phase, causing detectors to generate false alarms Abbott et al. (2018a). The presence of glitches causes incorrect sky positioning in GW data analysis, affects the tracking of electromagnetic (EM) counterparts Pankow et al. (2018), and introduces biases in the inference of cosmological parameters Macas et al. (2022); Zhu and Chen (2023).
@@
The interference caused by glitches in GW data cannot be ignored, and many efforts are being made to mitigate their disturbance Powell (2018); Soni et al. (2020); Cornish (2021); Talbot et al. (2021); Steltner et al. (2022); Udall and Davis (2023); Hourihane et al. (2022); Mohanty and Chowdhury (2023). The gating approach uses a windowing function to directly smooth the portion affected by the glitch for events whose time-frequency can be separated Usman et al. (2016). Gwsubtract Davis et al. (2019); Abbott et al. (2023b); Davis et al. (2022) relies on the auxiliary channel to remove glitches, but it has a limited impact on removing glitches that do not appear in auxiliary channels. BayesWave reconstructs the Gaussian noise, glitches, and GW signals using wavelet transforms and inter-correlations between detectors to remove glitches Cornish and Littenberg (2015); Littenberg et al. (2016); Cornish et al. (2021). However, this high-dimensional Bayesian analysis is extremely computationally intensive, and individual events cannot be analyzed in a timely and efficient manner at expected future event rates. Therefore, the development of an effective framework for rapidly mitigating the impacts of glitches and inferring GW parameters has become an urgent need.
@@
The development of Artificial Intelligence offers an excellent way to facilitate GW research by overcoming the problems of high computational costs and handling huge amounts of data Huerta et al. (2021). Currently, most state-of-the-art machine learning algorithms remove non-Gaussian noise components Biswas et al. (2013); Vajente et al. (2020); Ormiston et al. (2020); Merritt et al. (2021); Yu and Adhikari (2022); Ashton (2023); Bini et al. (2023) by fitting their parameters. Moreover, machine learning techniques Gabbard et al. (2018); Chua et al. (2019); Wang et al. (2020); Chen et al. (2020); Cabero et al. (2020); Green and Gair (2021); Gunny et al. (2022); Ruan et al. (2023a); Zhao et al. (2023); Dax et al. (2023); Ruan et al. (2023b); Yun et al. (2023); Ma et al. (2024); Zhao et al. (2024a) utilize the normalizing flow Williams et al. (2021); Ruhe et al. (2022); Langendorff et al. (2023); Du et al. (2024); Sun et al. (2024a) for the posterior distribution of source parameters Fan et al. (2019); Chua and Vallisneri (2020); Dax et al. (2021); Gabbard et al. (2022). While these methods can effectively enhance data quality Zhao et al. (2024b) in certain aspects, they cannot ensure that the noise remaining after deglitching consists solely of Gaussian noise. Moreover, other approaches have not involved the removal or subtraction of non-Gaussian noise Du et al. (2024); Sun and Li (2023).
@@
In our previous work, we employed the normalizing flow to perform parameter inference on GW signals contaminated by several common glitches Sun et al. (2024b). However, glitches come in different types with varying frequencies of occurrence, and not all types of glitches can be well modeled or have sufficient numbers for training. We are seeking a method that does not rely on modeling to remove glitches and enable rapid parameter inference. Recent studies have demonstrated the robustness of machine learning Wei and Huerta (2020); Chatterjee et al. (2021); Wang et al. (2024a, b); Xu et al. (2024). The robustness of machine learning methods can aid in performing parameter inference reasonably, even without explicitly modeling and removing glitches. It enables the effective mitigation of non-Gaussian noise interference while inferring GW parameters.
@@
In this work, our primary aim is to investigate the impact of noise transients on the parameter inference obtained from the normalizing flow network. We use Gaussian noise for training to infer signal parameters contaminated by non-Gaussian noise. Specifically, we utilize simulated sine-Gaussian glitches, encompassing a wide range of morphologies found in advanced detector noise transients Powell et al. (2015); Bose et al. (2016). We discuss the effect of the number of detectors on robust inference in the normalizing flow network. We find that our model provides a relatively better posterior distribution compared to the Bayesian inference method when dealing with glitch-contaminated GW data. In this work, we utilize the bilby dynesty Ashton et al. (2019); Speagle (2020) to implement Bayesian inference. We have also explored the impact of the GW signal intensity on the robustness of network inference. The results indicate that the SNR of the GW signal has varying degrees of impact on the network’s ability to perform source parameter inference. Furthermore, we consider the influence of glitches’ intensity on the robustness of source parameter inference when utilizing the normalizing flow network. Meanwhile, we conduct a detailed study on the impact of three key features of non-stationary noise Canton et al. (2014), on the robustness inference using the normalizing flow. We find that it is possible to determine the feasibility of obtaining reasonable results by assessing the robustness of the inference considering the characteristics of non-Gaussian noise.
@@
The results demonstrate that the normalizing flow network can provide reliable posterior parameter inference for GWs contaminated by unknown glitches without undergoing deglitching. When GW data are contaminated by unknown noise, our model is capable of delivering swift and reasonable parameter inference, albeit with a difference in magnitude. This provides a reasonable prior for subsequent more accurate processing. When extended to binary neutron star (BNS) mergers, this method is also expected to enable the rapid localization of EM counterparts.
@@
This paper is structured as follows. In Section II, the methodology used in the work is described in detail. Some of the work’s conclusions and discussions are presented in Section III. Finally, we present the main conclusions of our work in Section IV.
@@ II.1 Normalizing flow network architecture
The normalizing flow is a deep learning-based generative model for learning complex probability distributions Rezende and Mohamed (2015). Its core idea is to transform a simple probability distribution into a complex probability distribution through a series of reversible transformations. Neural posterior estimation Papamakarios and Murray (2016) involves training a neural network to approximate the transformation process of this complex posterior distribution, thereby achieving an approximate estimation of the posterior distribution. In the inference process, the normalizing flow generates samples by applying a series of reversible transformations to the base distribution. The normalizing flow network learns the parameters of these transformations to map the base distribution to a posterior distribution corresponding to the real data distribution.
The normalizing flow network consists of a series of coupling transformations, each of which is transformed according to the input data and the output of the previous layer of the neural network, allowing the predicted distribution to gradually approximate the posterior distribution. We also employed neural spline flow Durkan et al. (2019), which utilizes monotonic rational quadratic spline coupling transformations. Between coupled transformations, sample components were inverted to ensure that the transformation sequence specifically and completely transformed all components.
The normalizing flow Rezende and Mohamed (2015); Kingma et al. (2016); Chen et al. (2016); Papamakarios et al. (2017) can be expressed by the following formula
@@
Our architecture for the normalizing flow is quite similar to the one described in Ref. Sun et al. (2024b), except for data fusion. Our network architecture combines a one-dimensional ResNet-50 network He et al. (2016) for feature extraction with the normalizing flow network for estimating the parameters of the posterior distribution. The overall structure of the working principle is illustrated in Fig. 1. The data strain input to the network first undergoes the ResNet neural network, then the feature vectors extracted from the ResNet model. Finally, the feature vectors are processed by the normalizing flow to obtain the GW waveform as well as the parameter inference.
Unlike the parameters used in Ref. Sun et al. (2024b), the network parameters in this work are shown in Table 1. For networks corresponding to different numbers of detectors, their network parameters remain the same, except that the number of channels in the network corresponds to the number of detectors. We also experimented with model complexity reduction and regularization techniques to enhance the robustness of the network but found that they had minimal impact on the results.
During the training process, we utilize the AdamW optimizer Loshchilov and Hutter (2018) with a learning rate set to 0.0001, a batch size of 200, and a learning rate decay factor of 0.99. To ensure the validity of the normalizing flow network predictions and prevent overfitting, we regenerate 10,000 new data points at the beginning of each epoch and employ 50 cycles of early stopping. The training of this network was conducted on an NVIDIA RTX A6000 GPU with 48 GB of memory.
Figure 1: The workflow of the normalizing flow network. The input strain data contain simulated GW signals and Gaussian noise. The strain data is passed through an identical residual net (ResNet) model, and then the feature vectors are extracted from the ResNet model. Then it proceeds through the normalizing flow to obtain the parameter posterior distributions. The above steps are repeated by injecting glitches into the strain data as a test set.
@@ II.2 Data set
In terms of dataset preparation, the data used during the training phase consists of simulated Gaussian noise and simulated GW signals. The signals used are from the GW signals simulated by the SEOBNRv4_opt model Bohé et al. (2017), and they are selected in the range of SNR between 8 and 32. The noise used is a simulated Gaussian noise generated using Pycbc Biwer et al. (2019); Dal Canton et al. (2014) with a frequency range of $20-4096$ Hz, which represents the high-pass frequency and sampling frequency respectively. This also implies a Nyquist frequency of 2048 Hz. To obtain the noise sensitivity curves ${S_{n,k}(f)}$ for different detectors, we utilize the aLIGOZeroDetHighPower, AdvVirgo, and KAGRA methods from the pycbc.psd library Nitz and othes (2022); Sellers et al. (2022). After generating Gaussian noise, we inject a signal between 0.7 s and 0.9 s. The specific parameters distribution of the signal are shown in Table 2, and the simulated GW observation data are whitened. Finally, a 2-second dataset is generated.
@@
where $\bm{\tilde{h}}$ is the frequency domain GW waveform considering the detector network including $N$ independent detectors and can be written as $\bm{\tilde{h}}=\left[\tilde{h}_{1},\tilde{h}_{2},\cdot\cdot\cdot,\tilde{h}_{k},\cdot\cdot\cdot,\tilde{h}_{N}\right]$. Here $\tilde{h}_{k}$ is the frequency domain GW waveform of the $k$-th detector. The inner product is defined as Jin et al. (2024a)
@@
and ${S_{n,k}(f)}$ is the sensitivity curve function of the $k$-th detector Owen and Sathyaprakash (1999). In the above expression, $f_{\rm min}$ is the high-pass frequency, and $f_{\rm max}$ is the Nyquist frequency. Also, we calculate the glitch SNR in the same way.
@@
In the test dataset, we utilize simulated Gaussian noise and simulated GW signals, injecting simulated glitches into them, as illustrated in Fig. 2. The glitches we use are modeled using sine-Gaussian functions Blackburn et al. (2008), which cover most of the types of glitches currently observed by the LIGO, Virgo, and KAGRA networks. Note that when injecting glitches into detectors, we selected the LIGO Livingston detector due to its higher sensitivity and higher occurrence frequency of glitches, as random glitches typically occur in only one detector. We examine a sine-Gaussian glitch Powell et al. (2015), which is defined as
@@
where $\tau=Q/\sqrt{2}\pi f_{0}$ is the time duration. Here $h_{0}={\rm hrss}/\sqrt{\tau}$, where $\rm hrss$ is the root sum squared amplitude of the glitch, $Q$ is the dimensionless quality factor, $t_{0}$ and $f_{0}$ are the location of the sine-Gaussian in time and frequency respectively. We set $t_{0}=0.8$ s to simulate the scenario where the signal and glitch overlap (the underlying rationale is provided in Appendix A). The selected parameters range of glitches are shown in Table 3.